# Advancing Multimodal Reasoning: From Optimized Cold Start to Staged Reinforcement Learning

### v1

@@ 4.2 Stage 2: Textual RL
While MRL is indispensable for grounding reasoning across visual and textual inputs, intensive MRL training can inadvertently lead to a decline in purely textual capabilities, which we define as “textual capability decay”. To further elevate the model’s capacity for sophisticated abstract reasoning, we integrate a subsequent Textual Reinforcement Learning (TRL) phase. This stage aims to achieve both robust linguistic fluency and advanced reasoning. Linguistic fluency is restored and enhanced by fine-tuning on high-quality, text-only corpora focused on instruction-following and conversational abilities. Simultaneously, to foster advanced reasoning, the TRL phase exposes the model to complex, text-centric problem-solving tasks. This compels the model to refine and generalize intricate reasoning patterns, articulate multi-step thought processes with greater clarity, and master linguistic nuances essential for higher-order cognition.
For policy optimization during this TRL phase, we employ GRPO, augmented with our proposed PAD mechanism for efficient sample utilization. The reward function is multifaceted, designed to promote linguistic excellence, which encompasses both fluency and conciseness. Conciseness, in particular, is also encouraged by the efficient-length reward (Equation 6).
@@ 5.1 Experiments Setup
Datasets
The training of ReVisual-R1 follows our proposed three-stage methodology, utilizing carefully curated datasets for each phase. The cold-start phase employed approximately 40k pure text entries focused on establishing foundational language understanding. Subsequently, the Multimodal Reinforcement Learning (MRL) phase used approximately 26k diverse multimodal entries from our GRAMMAR dataset (see Section 3.2 for data curation details) to develop cross-modal reasoning. Finally, the text-based RL (TRL) phase consisted of approximately 30k text entries designed to refine nuanced understanding and generation capabilities.
@@
We evaluate ReVisual-R1 on a comprehensive suite of benchmarks, selected to test diverse reasoning skills. For visual-mathematical reasoning, we employed MathVerse zhang2024mathverse , MathVision wang2024measuring , WeMath qiao2024wemathdoeslargemultimodal , and DynaMath zou2025dynamathdynamicvisualbenchmark . Broader multimodal reasoning was assessed using MathVista lu2023mathvista and LogicVista xiao2024logicvistamultimodalllmlogical . Performance on challenging text-based mathematical reasoning was measured on AIME24/25 li2024numinamath and MATH-500 dataset_math , while general question answering was tested with GPQA rein2024gpqa . We report pass@1 accuracy for performance on the evaluated benchmarks except AIME. For the AIME24/25 benchmark, performance is measured using pass@32 accuracy.
@@
The performance of ReVisual-R1 is benchmarked against several categories of models, with detailed results presented in Table 2. These baselines include: (1) leading closed-source models such as doubao-1.5-vision-pro-32k guo2025seed1 , OpenAI-GPT-4o hurst2024gpt , Claude-3.7-Sonnet claude-3.7 , and Gemini-2.0-Flash team2023gemini ). (2) diverse open-source general-purpose MLLMs like the InternVL3-8B internvl2.5 , LLaVA-OneVision llavaov-li-arxiv-2024 , and the Qwen2.5-VL-7B bai2025qwen2 ; and (3) specialized open-source reasoning MLLMs, including VLAA-Thinker-7B chen2025sft , OpenVLThinker-7B openvlthinker , MMR1-Math-v0 MMR1-Math2025 , MM-Eureka mmeureka-meng-arxiv-2025 , and VL-Rethinker-7B wang2025vl .
Implementation
Our ReVisual-R1 model is based on the Qwen-2.5-VL-7B-Instruct model. Its training comprised three distinct stages. The process begins with a cold-start phase utilizing LLaMA Factory zheng2024llamafactory and pure text data to establish foundational language understanding. Following this, Multimodal Reinforcement Learning (MRL) is implemented using Easy R1 zheng2025easyr1 . In this stage, the GRPO Kullback-Leibler (KL) divergence constraint is omitted to encourage broader policy exploration. The final stage involves Text-based Reinforcement Learning (TRL), also conducted via Easy R1. During TRL, the vision tower is frozen to concentrate learning on textual reasoning, and a small KL penalty is incorporated alongside entropy annealing to enhance training stability. All experiments are conducted on a setup of 8 NVIDIA A100-80G GPUs. Detailed prompt settings and training hyperparameters are provided in the Appendix.
@@ 5.2 Main Results
As shown in Table 2, our model achieves state-of-the-art (SOTA) results on math-related benchmarks among open-source reasoning multimodal models and even beats some commercial large MLLM. These results prove the effectiveness of our method beneficial to the open-source community.
Specifically, ReVisual-R1 achieves an impressive average score of 53.1%, a significant improvement of +16.8 percentage points over the previous open-source SOTA average. Specifically, ReVisual-R1 secures the top position among open-source contenders in nine out of ten individual benchmarks: MathVerse (+5.4% $\Delta$), MathVision (+13.9% $\Delta$), DynaMath (+9.8% $\Delta$), WeMath (+0.2% $\Delta$), LogicVista (+9.6% $\Delta$), AIME24 (+44.6% $\Delta$), AIME25 (+15.4% $\Delta$), GPQA (+10.1% $\Delta$), and MATH500 (+23.4% $\Delta$). The most substantial gains are observed in the challenging AIME24, MATH500, and AIME25 benchmarks, underscoring ReVisual-R1’s advanced mathematical and inferential reasoning prowess. On MathVista, ReVisual-R1 achieves the second-best open-source score, narrowly trailing the leading open-source model by only -0.6%.
@@
Collectively, these results validate the efficacy of our proposed training method, including the structured three-stage curriculum and enhancements like Prioritized Advantage Distillation. The strong performance of ReVisual-R1 not only pushes the boundaries for open-source 7B MLLMs in complex reasoning but also provides a valuable contribution to the broader research community.
Table 2: Performance comparison of various MLLMs on diverse out-of-domain benchmarks. The best scores are bold; the second best are underlined (among open-source models). AIME24 and AIME25 results are averaged over eight independent inference runs to reduce score variance.
Multimodal Reasoning Benchmarks Textual Reasoning Benchmarks Model MathVerse MathVision MathVista DynaMath WeMath LogicVista AIME24 AIME25 GPQA MATH500 Avg. Close-Source doubao-1.5-vision-pro-32k 64.7 51.5 78.6 44.9 64.2 65.7 26.7 20.0 56.1 85.2 55.8 OpenAI-GPT-4o 40.6 31.1 59.9 34.5 42.9 64.4 9.3 8.3 49.9 74.6 41.6 Claude-3.7-Sonnet 52.0 41.3 66.8 39.7 58.2 49.3 20.0 13.3 61.1 80.4 48.2 Gemini-2.0-Flash 43.6 47.8 70.4 42.1 47.4 52.3 33.3 36.7 35.4 69.0 47.8 Open-Source General Models InternVL-3-8B 33.9 28.6 70.5 23.0 37.5 43.6 3.3 6.7 34.8 75.2 35.7 LLaVA-OV-7B 17.6 17.6 62.6 9.0 17.7 32.0 0.0 0.0 0.1 45.2 20.2 Qwen-2.5-VL-7B 38.7 26.6 69.1 12.6 24.5 35.6 10.0 6.7 32.8 67.2 32.4 Open-Source Reasoning Models OpenVLThinker-7B 38.1 23.0 65.3 16.8 35.2 44.5 5.0 1.7 28.3 51.0 30.9 MM-Eureka-Qwen-7B 45.4 28.1 72.6 23.0 21.8 46.3 6.7 3.3 34.3 66.6 31.5 MMR1-Math-v0 42.8 30.7 69.8 17.4 31.9 46.8 5.4 0.8 19.2 65.8 33.1 ThinkLite-7B-VL 42.9 24.6 71.6 16.5 41.8 42.7 8.8 27.9 24.8 61.4 36.3 VLAA-Thinker-7B 44.9 24.2 71.7 17.8 35.7 45.9 0.8 12.6 30.8 30.8 31.5 VL-Rethinker-7B 46.4 28.4 73.7 17.8 36.3 42.7 2.9 2.9 37.4 47.0 33.6 Our model ReVisual-R1 53.6 48.8 73.1 27.5 42.0 52.3 53.3 43.3 47.5 89.2 53.1 $\Delta$ (Ours–Open SoTA) +7.2 +18.1 -0.6 +4.5 +0.2 +5.5 +44.5 +15.4 +10.1 +23.4 +16.8
@@ 5.3.1 Ablation on SRO
Table 3: Ablation study of different training stage combinations applied to the ReVisual-R1 model, building upon a Cold Start. Best results per column are bold and second-best are underlined.
[…]
To validate our Staged Reinforcement Optimization (SRO) framework, we conducted ablation studies on different combinations of Multimodal RL (MRL) and Text-based RL (TRL) phases, all building upon our optimized text-centric cold-start (CS). This investigation, with results detailed in Table 3, aimed to empirically determine the most effective sequence for cultivating a balance between robust visual grounding and advanced textual proficiency.
The empirical evidence strongly supports our proposed CS + MRL + TRL (ReVisual-R1-MTR) sequence, which consistently yielded the highest average performance (49.6 Avg). This outcome affirms our core hypothesis: an initial MRL phase dedicated to establishing strong visual grounding, followed by a TRL phase to refine textual fluency and abstract reasoning, is crucial for developing superior multimodal capabilities without degrading the foundational cross-modal understanding.
In a more detailed analysis, the CS + MRL only model (47.7 Avg), while performing well on visually intensive tasks such as MathVista (71.9), did not reach the overall performance of the full MTR sequence. This suggests that MRL, while vital, can lead to a “textual capability decay”, which the subsequent TRL stage effectively mitigates. The alternative SRO ordering, CS + TRL + MRL (45.5 Avg), also proved less effective than our MTR approach. This finding indicates that establishing strong visual grounding before intensive textual refinement allows for more synergistic learning, where the TRL phase can enhance reasoning that is already connected across modalities.
In conclusion, these ablation results provide compelling justification for the MRL-then-TRL ordering within our SRO framework. This strategic sequencing first grounds the model multimodally and then sharpens its linguistic and abstract reasoning faculties, culminating in a more comprehensively capable and high-performing MLLM.
@@ 5.3.2 Ablation study on PAD
Table 4: Ablation results demonstrating the impact of Prioritized Advantage Distillation (PAD) and its core components. Performance metrics are reported for various mathematical reasoning benchmarks, averaged across datasets. Best results per column are bold and second-best are underlined.
[…]
We conduct ablation studies to evaluate Prioritized Advantage Distillation (PAD), examining its overall efficacy, the contribution of its components, and its sensitivity to key hyperparameters.
Figure 3: Training reward of different strategies: PAD-sampling, GRPO-Baseline, GRPO-Filter, and Random-Sampling. PAD consistently reaches higher accuracy and converges faster.
To assess PAD’s impact, its full implementation was compared against GRPO-Baseline, GRPO-Filter-only, and Random-Sampling strategies. Table 4 demonstrates that full PAD achieved superior performance on mathematical reasoning benchmarks, highlighting the importance of its core components: effective sample filtering and prioritized sub-sampling. Training dynamics (Figure 3) further corroborate PAD’s effectiveness, with its sampling strategy yielding higher reward accuracy and faster convergence, thereby enhancing learning efficiency.
5.3.3 Ablation on Efficient-Length Reward
Figure 4: Training dynamics comparing models with (purple lines, “w efficient-length”) and without (green lines, “w/o efficient-length”) the Efficient-Length Reward.
In this paper, in multimodal RL, we devise an Efficient-Length Reward. As depicted in Figure 4, the Efficient-Length Reward significantly impacts training. The regularized model maintained stable and higher reward accuracy (Fig. 4a) and consistently low entropy (Fig. 4b). In contrast, the baseline model suffered an accuracy decline and a dramatic entropy increase. Furthermore, the Efficient-Length Reward helped maintain a stable mean response length (Fig. 4c) and a low clip ratio (Fig. 4d), unlike the baseline, which exhibited uncontrolled growth in response length and a consequently higher clip ratio.
In summary, the Efficient-Length Reward is crucial for stabilizing training, preventing accuracy degradation, maintaining low model entropy, and controlling verbosity.

### v2

@@ 4.2 Stage 2: Textual RL
While MRL is indispensable for grounding reasoning across visual and textual inputs, intensive MRL training can inadvertently lead to a decline in purely textual capabilities. To further elevate the model’s capacity for sophisticated abstract reasoning, we integrate a subsequent Textual Reinforcement Learning (TRL) phase. This stage aims to achieve both robust linguistic fluency and advanced reasoning. Linguistic fluency is restored and enhanced by fine-tuning on high-quality, text-only corpora focused on instruction-following and conversational abilities. Simultaneously, to foster advanced reasoning, the TRL phase exposes the model to complex, text-centric problem-solving tasks. This compels the model to refine and generalize intricate reasoning patterns, articulate multi-step thought processes with greater clarity, and master linguistic nuances essential for higher-order cognition. For policy optimization during this TRL phase, we employ GRPO, augmented with our proposed PAD mechanism for efficient sample utilization.
@@ 5.1 Experiments Setup
To comprehensively evaluate our model’s performance, we selected a diverse suite of benchmarks targeting a wide range of reasoning abilities. The evaluation suite is organized by domain: (1) Multimodal mathematical reasoning, assessed using MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024a), MathVista (Lu et al., 2023a), DynaMath (Zou et al., 2025), and WeMath (Qiao et al., 2024a); (2) General multimodal reasoning, evaluated with LogicVista (Xiao et al., 2024), MMMU (Yue et al., 2024a), MMMU-Pro (Yue et al., 2024b), CMMMU (Zhang et al., 2024a), and MMStar Chen et al. (2024); and (3) Text-based reasoning, measured on challenging mathematical benchmarks like AIME24/25 (Li et al., 2024) and MATH-500 (Hendrycks et al., 2021), as well as general reasoning benchmarks such as GPQA (Rein et al., 2024) and MMLU Pro (Wang et al., 2024b). Performance is reported as pass@1 accuracy for all benchmarks, with the exception of AIME24/25, for which we use average@32 accuracy.
@@
As shown in Table 2, baselines include: (1) leading closed-source models doubao-1.5-vision-pro-32k (Guo et al., 2025b), OpenAI-GPT-4o (Hurst et al., 2024), Claude-3.7-Sonnet (Anthropic, 2024), Gemini-2.0-Flash (Gemini Team et al., 2023). (2) open-source general-purpose MLLM Qwen2.5-VL-7B (Bai et al., 2025); and 3) specialized open-source reasoning MLLMs VLAA-Thinker-7B (Chen et al., 2025a), OpenVLThinker-7B (Deng et al., 2025), MMR1-Math-v0 (Leng et al., 2025), MM-Eureka (Meng et al., 2025a), and VL-Rethinker-7B (Wang et al., 2025).
5.2 Implementation Details
Our ReVisual-R1 models are based on the Qwen-2.5-VL-7B-Instruct and Qwen-2.5-VL-3B-Instruct models. Its training comprised three distinct stages. The process begins with a cold-start phase utilizing LLaMA Factory (Zheng et al., 2024) and pure text data to establish foundational language understanding. Following this, Multimodal Reinforcement Learning (MRL) is implemented using Easy R1 (Zheng et al., 2025). In this stage, the GRPO Kullback-Leibler (KL) divergence constraint is omitted to encourage broader policy exploration. The final stage TRL is also conducted via Easy R1. During TRL, the vision tower is frozen to concentrate learning on textual reasoning, and a small KL penalty is incorporated alongside entropy annealing to enhance training stability. All experiments are conducted on a setup of 8 NVIDIA A100-80G GPUs. Detailed prompt settings and training hyperparameters are provided in the Appendix.
5.3 Training Datasets
The training of ReVisual-R1 follows our proposed three-stage methodology, utilizing carefully curated datasets for each phase. The cold-start phase employed approximately 283k pure text entries focused on establishing foundational language understanding. Subsequently, the Multimodal Reinforcement Learning (MRL) phase used approximately 26k diverse multimodal entries from our GRAMMAR dataset to develop multimodal reasoning. Finally, the TRL phase consists of approximately 30k text entries designed to refine nuanced understanding and generation capabilities.
Table 2: Performance comparison on diverse benchmarks. The best scores are bold; the second best are underlined (among open-source models). Scores in italics indicate that they are not reported in the original work and are obtained using the VLMEvalKit (Duan et al., 2024) for evaluation. AIME24 and AIME25 results are averaged over eight independent inference runs to reduce score variance. MathVerse-V, DynaMath-W and WeMath-S denotes the vision-only, worst, and strict settings, respectively. $\Delta$ (e.g., Ours-Open 3B Best) denotes the improvement margin of the corresponding ReVisual-R1 model over the best-performing open-source baseline model in the same scale across the respective column.
Multimodal Reasoning Benchmarks Textual Reasoning Benchmarks Model MathVerse-V MathVision MathVista DynaMath-W WeMath-S LogicVista AIME24 AIME25 GPQA MATH500 Avg. Closed‑Source doubao-1.5-vision-pro-32k 64.7 51.5 78.6 44.9 64.2 65.7 26.7 20.0 56.1 85.2 55.8 OpenAI‑GPT‑4o 40.6 31.1 59.9 34.5 42.9 64.4 9.3 8.3 49.9 74.6 41.6 Claude‑3.7‑Sonnet 52.0 41.3 66.8 39.7 58.2 49.3 20.0 13.3 61.1 80.4 48.2 Gemini‑2.0‑Flash 43.6 47.8 70.4 42.1 47.4 52.3 33.3 36.7 35.4 69.0 47.8 3B-scale MLLMs Qwen2.5-VL-3B 33.0 22.7 60.0 9.0 17.9 28.8 6.7 0.0 22.2 57.2 25.8 FAST-3B 44.0 26.8 63.8 15.4 23.3 35.4 6.7 3.3 25.3 55.4 29.9 VLAA-Thinker-3B 36.4 24.4 61.0 18.2 33.8 38.5 3.3 0.0 27.8 61.4 30.5 ReVisual-R1-3B 46.2 46.0 64.8 24.6 39.6 45.4 40.0 36.7 37.4 84.2 46.5 $\Delta$ (Ours-Open 3B Best) +2.2 +19.2 +1.0 +6.4 +5.8 +6.9 +33.3 +33.4 +9.6 +22.8 +16.0 7B-scale Models Qwen‑2.5‑VL‑7B 38.7 26.6 68.2 12.6 24.5 35.6 10.0 6.7 32.8 67.2 32.3 OpenVLThinker‑7B 38.1 25.9 72.3 16.8 35.2 44.5 5.0 1.7 28.3 51.0 31.9 MM-Eureka-Qwen-7B 45.4 26.9 73.0 23.0 21.8 46.3 6.7 3.3 34.3 66.6 34.7 MMR1-Math-v0 45.1 30.2 71.0 17.4 30.2 50.8 5.4 0.8 19.2 65.8 33.6 ThinkLite‑7B‑VL 42.9 24.6 71.6 16.5 41.8 42.7 8.8 27.9 24.8 61.4 36.3 VLAA-Thinker-7B 48.2 26.4 68.0 22.4 41.5 48.5 0.8 12.6 30.8 30.8 33.0 VL-Rethinker-7B 46.4 28.4 73.7 17.8 36.3 42.7 2.9 2.9 37.4 47.0 33.6 ReVisual-R1-7B 53.6 48.8 73.1 27.5 42.0 52.3 53.3 43.3 47.5 89.2 53.1 $\Delta$ (Ours-Open 7B Best) +5.4 +18.6 -0.6 +4.5 +0.2 +1.5 +43.3 +15.4 +10.1 +22.0 +16.8
@@ 5.4 Main Results
As shown in Table 2, our model demonstrates superior performance on math-related benchmarks compared to other open-source reasoning models and even outperforms some commercial MLLMs.
Specifically, on the 7B model scale, ReVisual-R1-7B achieves an impressive average score of 53.1%, a significant improvement of +16.8 percentage points over the baselines in Tabel 2 on textual and multimodal benchmarks. Notably, ReVisual-R1 secures the top position among open-source contenders in nine out of ten individual benchmarks: MathVerse (+5.4% $\Delta$), MathVision (+18.6% $\Delta$), DynaMath (+4.5% $\Delta$), WeMath (+0.2% $\Delta$), LogicVista (+1.5% $\Delta$), AIME24 (+43.3% $\Delta$), AIME25 (+15.4% $\Delta$), GPQA (+10.1% $\Delta$), and MATH500 (+22.0% $\Delta$). All these results present the superior performance of our Revisual-R1 on both multimodal benchmarks as well as textual benchmarks.
@@
To further demonstrate the effectiveness of our training framework, we conduct experiments on 3B model scale, leading to Revisual-R1-3B. As demonstrated in Table 2, our model also demonstrates superior performance on this model size. Specifically, our model outperforms VLAA-Thinker-3B across all benchmarks and obtains an average 16.0% improvement. These results further verify the generalization of our Revisual-R1.
Collectively, these results validate the efficacy of our proposed training method, including the structured three-stage curriculum and enhancements like Prioritized Advantage Distillation.
Figure 4: Performance comparison on general textual and multimodal benchmarks.
@@ 5.5.1 Ablation on SRO
Table 3: Ablation study of different training stage combinations applied to the ReVisual-R1 model, building upon a Cold Start. Best results per column are bold and second-best are underlined. Mixed-RL denotes that the model is jointly optimized with both MRL and TRL objectives in a mixed training stage.
Training Strategies MathVerse-V MathVision MathVista DynaMath-W WeMath-S LogicVista Avg Cold Start (CS) only 51.9 $47.9$ $70.5$ 26.5 $35.8$ $50.1$ $47.1$ CS + MRL $50.9$ $47.6$ 71.9 $25.7$ 38.8 51.2 $47.7$ CS + TRL $47.3$ $47.3$ $71.0$ $25.2$ $33.7$ $44.7$ $44.9$ CS + MRL + TRL 53.6 48.8 73.1 27.5 42.0 52.3 $49.6$ CS + TRL + MRL $47.5$ 48.0 $70.3$ $24.2$ $35.0$ $48.2$ $45.5$ CS + Mixed-RL $49.3$ $48.2$ $72.1$ $25.7$ $38.8$ $51.2$ $47.6$
To validate the effectiveness of our Staged Reinforcement Optimization (SRO) framework, we conduct ablation studies on different combinations of Multimodal RL (MRL) and Textual RL (TRL) phases, all building upon our optimized text-centric cold-start (CS). As shown in Table 3, the empirical results verify that our proposed CS + MRL + TRL (ReVisual-R1-MTR) sequence consistently yields the highest average performance (49.6 Avg). This outcome affirms our core hypothesis: an initial MRL phase establishes strong visual grounding, followed by a TRL phase to refine textual fluency and abstract reasoning, is crucial for developing superior multimodal capabilities.
In a more detailed analysis, the CS + MRL only model (47.7 Avg), while performing well on visually intensive tasks such as MathVista (71.9), does not reach the overall performance of the full MTR sequence. This further highlights the importance of the subsequent TRL stage. The alternative SRO ordering, CS + TRL + MRL (45.5 Avg), also proved less effective than our MTR approach. This finding indicates that establishing strong visual grounding before intensive textual refinement allows for more synergistic learning.
In addition, we perform a mixed-RL strategy to investigate the necessity of our staged reinforcement learning. In this setting, we train the cold-start model on a mixture of multimodal and textual data. However, as shown in Table 3, we observe that this mixed setting can only get an average 47.6 score, which is significantly lower than the performance of our staged RL. It shows that our MRL-then-TRL ordering within our SRO framework is a more effective strategy than simultaneously training on mixed-modality data in the RLVR stage.
@@ 5.5.2 Ablation study on PAD
Table 4: Ablation results demonstrating the impact of Prioritized Advantage Distillation (PAD) and its core components. Best results per column are bold and second-best are underlined.
Model Configuration Strategy MathVerse MathVision MathVista DynaMath WeMath LogicVista Avg ReVisual-R1 (CS + MRL) PAD 50.9 47.6 71.9 25.7 38.8 51.2 $47.7$ w/o PAD Components: - Full PAD (Baseline) GRPO-Baseline $47.6$ $45.8$ $68.8$ $25.2$ $34.8$ $48.6$ $45.1$ - No Prioritized Sub-sampling GRPO-Filter $47.7$ 46.7 71.2 $25.5$ $35.1$ 49.7 $46$ - No Effective Sample Filtering Random-Sampling 47.9 $46.4$ $70.7$ 26.1 37.1 $49.3$ $46.2$ Other Strategy DAPO 48.3 $46.3$ $69.2$ 25.4 38.3 $49.2$ $46.1$
In this section, we conduct ablation studies to verify the effectiveness of our proposed Prioritized Advantage Distillation (PAD), examining its overall efficacy, the contribution of its components, and its sensitivity to key hyperparameters.
To assess PAD’s impact, its full implementation is compared against GRPO-Baseline, GRPO-Filter-only, and Random-Sampling strategies. Table 4 demonstrates that full PAD achieved superior performance on mathematical reasoning benchmarks, highlighting the importance of its core components: effective sample filtering and prioritized sub-sampling. Meanwhile, we compare our method with DAPO and observe that our PAD can also present significantly better performance than DAPO. To further demonstrate the effectiveness of our Training dynamics (Figure 5) further corroborate PAD’s effectiveness, with its sampling strategy yielding higher reward accuracy and faster convergence.
Figure 5: Ablation of training dynamics of our PAD.
5.5.3 Generalization of Revisual-R1
To further evaluate the generalization capacity of our method, we test the model on several knowledge-intensive benchmarks, including MMMU, MMMU-Pro, and CMMMU, as well as on general-purpose multimodal perception (MMStar) and textual understanding (MMLU). The results, presented in Figure 4, show that our method consistently performs superior across these benchmarks. The more detailed performance is shown in Table 6. It further indicates that the generality of our approach on general tasks and provides further evidence that textual reasoning capabilities can benefit on general multimodal and textual tasks.